what is the focus of the parabola with the equation x^2=48y
Answer by stanbon(75887)

what is the focus of the parabola with the equation x^2=48y
----
Form: (x-0)^2 = 4p(y-0)
----
4p = 48
p = 12
-------
Vertex is at (0,0)
Focus is at (0,0+p) = (0,12)
------------
Cheers,
Stan H.